OpenStudy (anonymous):

one number is 2 more than twice another. the sum of the number is 14. find the numbers?

OpenStudy (anonymous):

n=2m n+m=14

OpenStudy (anonymous):

that dont work

OpenStudy (anonymous):

m = 2n + 2 2n + 2 + n = 3n + 2 = 14 3n = 12 n = 4 m = 10

OpenStudy (anonymous):

how did you get 2n+2+n=3n+2=14 m=2n+n then what

OpenStudy (anonymous):

m = 2n + 2 because 2 more than twice another

OpenStudy (anonymous):

got that

OpenStudy (anonymous):

next step?

OpenStudy (anonymous):

sum of m and n from the previous step, you get m = 2n + 2 m + n = 2n + 2 + n

OpenStudy (anonymous):

ok ty